# Spatial-Spectral Morphological Mamba for Hyperspectral Image Classification

### v1

@@ II-B Token Enhancement and Multi-head Attention
This module includes a gating mechanism that works with spatial and spectral tokens, using the central region of the HSI patch for contextual information [49]. This allows MorpMamba to dynamically adjust the importance of different tokens, improving its feature extraction capabilities. The enhancement of tokens occurs in the following manner:
@@
In this context, $\mathbf{W}_{spectral}$ and $\mathbf{W}_{spatial}$ represent the learned weight matrices, while $\sigma$ denotes the sigmoid function, as illustrated in Figure 2. This gating mechanism allows the model to dynamically enhance the importance of spatial and spectral features based on contextual information from the central region of the HSI sample.
Figure 2: Spatial-Spectral Token Enhancement module adopted to refine and enhance the spatial and spectral features.
To further enhance the model’s learning capabilities, we employ a multi-head attention mechanism. This allows the model to attend to different segments of the input tokens simultaneously, as illustrated in Figure 3. By projecting the input tokens into multiple attention heads, the model can capture complex dependencies within the HSI data, which is crucial for effective feature extraction. For each attention head $i$, the computations are defined as follows:
Figure 3: Multi-Head Self-Attention Module adopted to process the enhanced spatial and spectral features.
@@
where $\mathbf{W}^{Q}_{i}$, $\mathbf{W}^{K}_{i}$, and $\mathbf{W}^{V}_{i}$ are learned weight matrices associated with the query, key, and value projections, respectively. The attention scores for each head $i$ are computed using scaled dot-product attention:
@@
Each attention head works with key vectors of dimension $d_{k}$. The output for each head, denoted by $O_{i}$, is calculated as $O_{i}=A_{i}V_{i}$. These individual outputs are then combined using concatenation: $O=\operatorname*{\parallel}(O_{1},O_{2},\dots,O_{h})$. This combined output, $O$, represents a fused feature representation. It captures diverse and refined features from both the spatial and spectral domains of the HSI data. This enrichment enhances the model’s ability to extract meaningful information.
@@ II-C State Space Model (SSM)
MorpMamba captures complex relationships within HSI data by combining several techniques: spectral-spatial token generation, token enhancement, multi-head attention, and an SSM. It starts with a sequence of improved tokens, denoted by $O=(E_{1},E_{2},E_{3},\dots,E_{T})$. The state transition, which captures these intricate dependencies, is then computed as:
$h_{t}=\text{ReLU}(W_{\text{transition}}h_{t-1}+W_{\text{update}}E_{t})$ (14)
where $W_{transition}$ and $W_{update}$ denote learned weights, and $h_{t}$ represents the hidden state at time $t$. To obtain the final output for classification, a linear classifier is applied to $h_{t}$:
@@
where $|W_{\text{classifier}}|_{2}^{2}$ represents the squared $l_{2}^{2}$ norm of the weights $W_{\text{classifier}}$. $\lambda=0.01$ is the regularization coefficient, controlling the strength of regularization and $\sigma$ denotes the sigmoid function used for classification. $l_{2}^{2}$ regularization penalizes large values in $W_{\text{classifier}}$, encouraging the model to prefer smaller and simpler models, which help prevent overfitting. This approach enables MorpMamba to effectively leverage both spatial and spectral information, enhancing its ability to classify HSI data by combining complementary insights from these modalities.
@@ III Experimental Datasets
To assess MorpMamba’s performance, we evaluated it on several publicly available HSI datasets. These datasets include WHU-Hi-LongKou, Pavia University (PU), Pavia Centre (PC), Salinas (SA), and the University of Houston (UH). Table I summarizes the key characteristics of each dataset used in our experiments and true maps are presented in Figure 4.
@@
Figure 6: Training Time of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets. The training ratio and patch size have a strong influence on computational time, whereas the head size within multi-head self-attention and the kernel size within morphological operations do not significantly affect the computational load. This demonstrates that the Mamba model maintains a linear computational load even after incorporating multi-head self-attention and morphological operations. However, these additions significantly improve performance, as shown in the subsequent sections.
@@ IV Parameter Tuning
Figure 8: The predicted ground truth maps for LK datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
@@
Figure 9: The predicted ground truth maps for PC datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
@@
Figure 10: The predicted ground truth maps for PU datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved slightly higher accuracy, ranging from 0.32% to 1.12%, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
@@
Figure 11: The predicted ground truth maps for SA datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
@@
Figure 12: The predicted ground truth maps for UH datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices. In contrast, the MorpMamba model offers significantly better parameter efficiency and overall performance compared to all competing methods.
@@ V Comparison with SOTA Methods and Discussion
In this section, we present a detailed comparative analysis of the proposed MorpMamba model against several SOTA HSIC models. The comparison is based on key performance metrics: Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). Additionally, we assess computational efficiency by evaluating the number of parameters for each model. Table II provides a comprehensive overview of these metrics across five prominent HS datasets: the UH, LK, PU, PC, and SA. Whereas, Figures 8, 9, 10, 11, and 12 present the predicted ground truth maps for all the competing methods, including the proposed MorpMamba model. OA represents the percentage of correctly classified samples out of the total samples, AA is the mean accuracy across all classes, and $\kappa$ is a statistical measure of inter-rater agreement adjusted for chance agreement.
@@
The 2D CNN architecture initiates with four sequential convolutional layers with 8, 16, 32, and 64 filters with a $3\times 3$ kernel size, using the same padding and ReLU activation. The convolutional layers are succeeded by a flattening layer that transforms the 2D feature maps into a 1D vector, which is then processed by two dense layers. Two dense layers consist of 256 and 128 units with ReLU activation and are followed by a Dropout layer with a dropout rate of 0.4 to mitigate overfitting. The network concludes with an output layer that uses the softmax activation function to provide classification probabilities, where the number of units corresponds to the number of classes.
The 3D CNN architecture starts with a series of four 3D convolutional layers. The first layer applies 8 filters with a kernel size of $(3,3,7)$, using ’same’ padding and ReLU activation. The second layer increases the filter count to 16, with a kernel size of $(3,3,5)$ and the same padding and activation functions. The third layer utilizes 32 filters with a $(3,3,3)$ kernel size, followed by a fourth layer with 64 filters and the same kernel dimensions. After the convolutional layers, the feature maps are flattened into a 1D vector. This vector is then processed through two fully connected (dense) layers: the first with 256 units and ReLU activation, and the second with 128 units, also using ReLU activation. The network concludes with a dense output layer using the softmax activation function to generate classification probabilities.
The HybCNN model integrates both 3D and 2D convolutional layers to leverage the strengths of spatial and spectral feature extraction. The model begins with three successive 3D convolutional input layers: the first applies 8 filters with a kernel size of $(3,3,7)$, the second applies 16 filters with a kernel size of $(3,3,5)$, and the third applies 32 filters with a kernel size of $(3,3,3)$, all using ’same’ padding and ReLU activation. After the 3D convolutions, the output is reshaped from a 5D tensor to a 4D tensor, combining the depth dimensions into a single dimension. This reshaped tensor is then processed by a 2D convolutional layer with 64 filters and a kernel size of $(3,3)$, followed by a flattening layer. The flattened output is passed through two dense layers: the first with 256 units and ReLU activation, followed by a Dropout layer with a rate of 0.4, and the second with 128 units and ReLU activation, also followed by a Dropout layer with a rate of 0.4. The model concludes with a dense output layer employing the softmax activation function to produce classification probabilities.
The 2D IN model captures multi-scale features from the input patch and comprises three distinct blocks of convolutional operations. In Block 1, the network applies a sequence of three convolutional layers: the first with 30 filters and a $(1,1)$ kernel size, the second with 20 filters and a $(1,1)$ kernel size, and the third with 10 filters and a $(3,3)$ kernel size, all using ’same’ padding and ReLU activation. Block 2 employs a different configuration, starting with a $(1,1)$ convolutional layer with 40 filters, followed by a $(1,1)$ convolutional layer with 20 filters, and concluding with a $(5,5)$ convolutional layer with 10 filters. Block 3 initiates with a max pooling layer of size $(3,3)$, followed by two convolutional layers: the first with 20 filters and a $(1,1)$ kernel size, and the second with 10 filters and a $(1,1)$ kernel size. The outputs from these three blocks are concatenated along the channel dimension, and the combined feature maps are processed by an additional convolutional layer with 128 filters and a $(1,1)$ kernel size. The network concludes with a classification module that includes a flattening layer, followed by three dense layers with 1200, 600, and 150 units, respectively, all using ReLU activation. The final output layer employs softmax activation to generate class probabilities.
The 3D IN model is a 3D-CNN designed to capture spatiotemporal features from volumetric input data effectively. The network architecture includes three distinct convolutional blocks. In Block 1, the model starts with a 3D convolutional layer featuring 30 filters with a kernel size of $(5,5,7)$, followed by another 3D convolutional layer with 20 filters and a kernel size of $(3,3,5)$, and concludes with a third 3D convolutional layer with 10 filters and a kernel size of $(3,3,3)$. Block 2 mirrors this configuration but with increased filter counts: the first layer uses 40 filters, the second uses 20 filters, and the third also uses 10 filters. Block 3 further enhances feature extraction with 60 filters in the first layer, and 30 filters in the second, and maintains 10 filters in the final layer, each using the same kernel sizes as in previous blocks. The outputs from these blocks are concatenated along the channel dimension, and the combined features are processed by an additional 3D convolutional layer with 128 filters and a $(1,1,1)$ kernel size. The network then transitions to a classification module, which includes a flattening layer followed by three dense layers. The first dense layer, with 512 units, is followed by a Dropout layer with a rate of 0.4. The subsequent dense layer, with 128 units, is also followed by a Dropout layer, and the final dense layer, with 64 units, is followed by another Dropout layer. The output layer employs softmax activation to produce class probabilities.
The HybIN model is a hybrid architecture combining both 3D and 2D convolutional layers to leverage the strengths of each for feature extraction from volumetric input data. This architecture also begins with three distinct 3D convolutional blocks. In Block 1, the model uses a 3D convolutional layer with 30 filters and a kernel size of $(5,5,7)$, followed by layers with 20 and 10 filters and kernel sizes of $(3,3,5)$ and $(3,3,3)$, respectively. After extracting features with 3D convolutions, the output is reshaped into a 2D structure, applying max pooling and subsequent 2D convolutions with 16, 32, and 64 filters, each with a $(1,1)$ kernel size. Block 2 follows a similar process but with different filter counts: starting with 40 filters, then 20, and ending with 10, each using the same kernel sizes as Block 1. After reshaping the output to 2D, max pooling and 2D convolutions with 16, 32, and 64 filters are applied. Block 3 increases the depth of the convolutional layers, using 60 filters in the first layer, 30 in the second, and 10 in the third, followed by 2D reshaping, max pooling, and 2D convolutions with 16, 32, and 64 filters. The outputs from all three blocks are concatenated along the channel dimension and processed by an additional 2D convolutional layer with 128 filters and a $(1,1)$ kernel size. The classification module includes flattening followed by three dense layers with 512, 128, and 64 units, respectively, each with ReLU activation and Dropout for regularization. The final layer uses softmax activation to produce class probabilities.
The Hybrid_ViT model combines 3D convolutional layers with Transformer layers to effectively capture both spatial and spectral features from volumetric input data. The model first applies a patch embedding layer using a 3D convolution with an embedding dimension of 64. This convolutional layer, with a kernel size of $(1,S,S)$, reduces the spatial dimensions and creates patch embeddings. A learnable class token is concatenated with these patch embeddings, and positional encodings are added to provide spatial context. The core of the model consists of several Transformer encoder layers, each with a configurable number of heads and depth. Each Transformer layer processes the patch embeddings to capture complex dependencies and relationships within the data. In addition to self-attention, the model incorporates a cross-attention mechanism between the class token and patch embeddings to enhance feature extraction. This cross-attention mechanism is implemented by computing the attention scores between the class token and patch embeddings, followed by a weighted aggregation of the patch embeddings based on these scores. Finally, the model concatenates the output of the Transformer layers with the cross-attention output and processes it through several dense layers with ReLU activation. The final output layer, using a softmax activation function, produces class probabilities based on the number of classes.
The hir_transformer models integrate pyramid-based 3D convolutional layers with Transformer-based mechanisms to leverage multi-scale features for enhanced volumetric data analysis. The model starts with a 3D input of shape $(S,S,B,1)$ and further constructs a pyramid of convolutional layers with varying scales, where each level of the pyramid applies 3D convolutions with increasing kernel sizes and scales. The input is scaled down for each pyramid level by a factor of $2^{l}$, where $l$ is the pyramid level. The output from each level, consisting of feature maps extracted at different scales, is then concatenated along the channel dimension to capture multi-scale information. The model extends this architecture by incorporating Transformer layers. It begins with the output from the pyramid_3d_cnn, which provides multi-scale features. The model applies a ReLU activation and processes the feature maps through several Transformer layers. Each Transformer layer consists of multi-head self-attention followed by a convolutional block, which includes 3D convolutions with kernel sizes designed to maintain spatial dimensions. The Transformer mechanism includes layer normalization and residual connections to ensure stable training and effective feature learning. Regularization techniques are employed to prevent overfitting. The feature maps are flattened and passed through dense layers with ReLU activation and L2 regularization, followed by a final softmax layer to produce class probabilities.
@@
To ensure a fair comparison across all methods in our experiments, we employed a consistent setup. We used a patch size of 4 and selected 15 bands for the analysis. The data was split with a training ratio of 40%, a validation ratio of 20%, and a test ratio of 40%. The models were trained over 50 epochs with a batch size of 256.
@@
We analyzed the performance of the MorpMamba model across various HS datasets, detailed in Section III. For the Uh dataset, MorpMamba achieves an OA of 98.28%, an AA of 97.91%, and a $\kappa$ of 98.14. Although the 3D CNN model shows slightly higher OA (99.01%) and AA (98.81%), MorpMamba remains competitive with significantly fewer parameters (67,013 compared to 4,042,751). On the LK dataset, MorpMamba attains an OA of 99.70%, an AA of 99.25%, and a $\kappa$ of 99.61, closely matching the top-performing models like the 3D CNN (OA of 99.81%) while excelling in computational efficiency with substantially fewer parameters (66,239 compared to 4,041,977). For the PU dataset, MorpMamba achieves an OA of 97.67%, an AA of 96.93%, and a $\kappa$ of 96.91, maintaining competitive performance and efficiency with a lower parameter count (66,239) compared to the 3D CNN model (OA of 98.70%). In the PC dataset, MorpMamba records an OA of 99.71%, an AA of 98.85%, and a $\kappa$ of 99.59, comparable to the 3D CNN model (OA of 99.87%) but with fewer parameters (66,239). Lastly, on the SA dataset, MorpMamba achieves an OA of 98.52%, an AA of 99.25%, and a $\kappa$ of 98.35. Although the 3D CNN remains slightly superior in OA (98.86%), MorpMamba’s performance is noteworthy given its reduced parameter count (67,142).
@@
The MorpMamba model demonstrates superior computational efficiency compared to Transformer-based models, which suffer from quadratic complexity with sequence length. By maintaining linear complexity, MorpMamba ensures scalability to larger datasets while providing competitive accuracy. Additionally, the use of morphological operations in MorpMamba enhances robustness and stability, effectively reducing noise and highlighting structural features in HS data. This approach mitigates challenges associated with high-dimensional data, as evidenced by consistent performance across diverse datasets. The comparative analysis underscores the strengths of the MorpMamba model in HSIC, consistently delivering high accuracy and efficiency, and outperforming or matching SOTA models while maintaining a lower computational footprint. These results validate the effectiveness of integrating morphological operations with the Mamba architecture, establishing MorpMamba as a leading model for HSIC.

### v2

@@ II-B Token Enhancement and Multi-head Attention
This module includes a gating mechanism that works with spatial and spectral tokens, using the central region of the HSI patch to obtain contextual information [39]. Additionally, this allows MorpMamba to dynamically adjust the importance of different tokens, improving its feature extraction capabilities. The enhancement of tokens occurs in the following manner:
@@
In this context, $\mathbf{W}_{spectral}$ and $\mathbf{W}_{spatial}$ represent the learned weight matrices, while $\sigma$ denotes the sigmoid function, as illustrated in Figure 2. This gating mechanism allows the model to dynamically enhance the importance of spatial and spectral features based on contextual information extracted from the central region of the HSI sample.
Figure 2: Spatial-Spectral token enhancement module adopted to refine and enhancement of the extracted spatial and spectral features.
We employ a multi-head self-attention mechanism to further enhance the model’s learning capabilities. This allows the model to attend to different segments of the input tokens simultaneously, as illustrated in Figure 3. By projecting the input tokens into multiple attention heads, the model can capture complex dependencies within the HSI data, which is crucial for effective feature extraction during the model training. For each attention head $i$, the computations are defined as follows:
Figure 3: Multi-head self-attention module adapted to interact with the enhanced spatial and spectral features.
@@
where $\mathbf{W}^{Q}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{k}}$, $\mathbf{W}^{K}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{k}}$, and $\mathbf{W}^{V}_{i}\in\mathbb{R}^{d_{\text{input}}\times d_{v}}$ are learned weight matrices associated with the query, key, and value projections, respectively, where for each attention head, $K_{i}\in\mathbb{R}^{N\times d_{k}}$, $Q_{i}\in\mathbb{R}^{N\times d_{k}}$, $V_{i}\in\mathbb{R}^{N\times d_{v}}$ be the dimension of key, query, and value. The attention scores for each head $i$ are computed using scaled dot-product between $\mathbf{Q_{i}}$ and $\mathbf{K_{i}}$:
@@
Each attention head works with key vectors of dimension $d_{k}$. The output for each head, denoted by $O_{i}$, is calculated as $O_{i}=A_{i}V_{i}$. These individual outputs are then combined using concatenation operations $O=\operatorname*{\parallel}(O_{1},O_{2},\dots,O_{h})$. This combined output $O$, represents a fused feature representation which captures diverse and refined features from both the spatial and spectral domains of the HSI data. This enrichment enhances the model’s ability to extract meaningful information in the context of classification.
@@ II-C State Space Model (SSM)
MorpMamba captures complex and intrinsic relationships within HSI data by combining techniques in several stages: spectral-spatial token generation, token enhancement, multi-head self-attention, and an SSM. It starts with a sequence of enhanced tokens, denoted by $O=(E_{1},E_{2},E_{3},\dots,E_{T})$. The state transition, which captures these intricate dependencies, is then computed as:
$h_{t}=\text{ReLU}(W_{\text{trans}}h_{t-1}+W_{\text{update}}E_{t})$ (14)
where $W_{trans}$ and $W_{update}$ denote learned weights, and $h_{t}$ represents the hidden state at time $t$. To obtain the final output for classification, a linear classifier is applied to $h_{t}$:
@@
where $|W_{\text{classifier}}|_{2}^{2}$ represents the squared $l_{2}^{2}$ norm of the weights $W_{\text{classifier}}$. $\lambda=0.01$ is the regularization coefficient, controlling the strength of regularization and $\sigma$ denotes the sigmoid function used for classification. $l_{2}^{2}$ regularization penalizes large values in $W_{\text{classifier}}$, encouraging the model to prefer smaller and simpler models, which help prevent overfitting. This approach enables MorpMamba to effectively leverage both spatial and spectral information, enhancing its ability to boost the classification performance by combining complementary insights from these modalities.
@@ III Experimental Datasets
To assess MorpMamba’s performance, we evaluated it on several publicly available HSI datasets. These datasets include WHU-Hi-LongKou, Pavia University (PU), Pavia Centre (PC), Salinas (SA), and the University of Houston (UH). Table I summarizes the key characteristics of each dataset used in our experiments.
@@
Figure 5: Training Time of MorpMamba across different training data ratios, patch sizes, number of heads, and kernel sizes over 50 epochs on WHU-Hi-LongKou, Pavia Centre, Pavia University, Salinas, and University of Houston datasets. The training ratio and patch size have a strong influence on computational time, whereas the head size within multi-head self-attention and the kernel size within morphological operations do not significantly affect the computational load.
@@ IV Parameter Tuning
Figure 7: The predicted ground truth maps for LK datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
@@
Figure 8: The predicted ground truth maps for PC datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
@@
Figure 9: The predicted ground truth maps for PU datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved slightly higher accuracy, ranging from 0.32% to 1.12%, their parameter efficiency remains constrained and too high for deployment on resource-limited devices as compared to MorpMamba.
@@
Figure 10: The predicted ground truth maps for SA datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices.
@@
Figure 11: The predicted ground truth maps for UH datasets using various SOTA methods, alongside the proposed MorpMamba model, are presented. While most competing methods achieved equivalent accuracy, their parameter efficiency remains constrained and too high for deployment on resource-limited devices.
@@ V Comparison with SOTA Methods and Discussion
In this section, we present a detailed comparative analysis of the proposed MorpMamba model against several SOTA HSIC models. The comparison is based on key performance metrics: Overall Accuracy (OA), Average Accuracy (AA), and the Kappa coefficient ($\kappa$). Additionally, we assess computational efficiency by evaluating the number of parameters for each model. Table II provides a comprehensive overview of these metrics across five prominent HSI datasets: the UH, LK, PU, PC, and SA. Whereas, Figures 7, 8, 9, 10, and 11 present the predicted ground truth maps for all the competing methods, including the proposed MorpMamba model. OA represents the percentage of correctly classified samples out of the total samples, AA is the mean accuracy across all classes, and $\kappa$ is a statistical measure of inter-rater agreement adjusted for chance agreement.
@@
To ensure a fair comparison across all methods in our experiments, we employed a consistent setup. We used a patch size of 4 and selected 15 bands for the analysis. The data was split with a training ratio of 40%, a validation ratio of 20%, and a test ratio of 40%. The models were trained over 50 epochs with a batch size of 256, Adam optimizer, and learning rate set to 0.001.
@@
We analyzed the performance of the MorpMamba model across various HSI datasets, detailed in Section III. For the UH dataset, MorpMamba achieves an OA of 98.28%, an AA of 97.91%, and a $\kappa$ of 98.14. Although the 3D CNN model shows slightly higher OA (99.01%) and AA (98.81%), MorpMamba remains competitive with significantly fewer parameters (67,013 compared to 4,042,751). On the LK dataset, MorpMamba attains an OA of 99.70%, an AA of 99.25%, and a $\kappa$ of 99.61, closely matching the top-performing models like the 3D CNN (OA of 99.81%) while excelling in computational efficiency with substantially fewer parameters (66,239 compared to 4,041,977). For the PU dataset, MorpMamba achieves an OA of 97.67%, an AA of 96.93%, and a $\kappa$ of 96.91, maintaining competitive performance and efficiency with a lower parameter count (66,239) compared to the 3D CNN model (OA of 98.70%). In the PC dataset, MorpMamba records an OA of 99.71%, an AA of 98.85%, and a $\kappa$ of 99.59, comparable to the 3D CNN model (OA of 99.87%) but with fewer parameters (66,239). Lastly, on the SA dataset, MorpMamba achieves an OA of 98.52%, an AA of 99.25%, and a $\kappa$ of 98.35. Although the 3D CNN remains slightly superior in OA (98.86%), MorpMamba’s performance is noteworthy given its reduced parameter count (67,142).
@@
The MorpMamba model demonstrates superior computational efficiency compared to Transformer-based models, which suffer from quadratic complexity with sequence length. By maintaining linear complexity, MorpMamba ensures scalability to larger datasets while providing competitive accuracy. Additionally, the use of morphological operations in MorpMamba enhances robustness and stability, effectively reducing noise and highlighting structural features in HSI data. This approach mitigates challenges associated with high-dimensional data, as evidenced by consistent performance across diverse datasets. The comparative analysis underscores the strengths of the MorpMamba model in HSIC, consistently delivering high accuracy and efficiency, and outperforming or matching SOTA models while maintaining a lower computational footprint. These results validate the effectiveness of integrating morphological operations with the Mamba architecture, establishing MorpMamba as a leading model for HSIC.